# Lost in Cultural Translation: Do LLMs Struggle with Math Across Cultural Contexts?

### v1

@@ 3 Countries Selection
  • Multidimensional Poverty Index (MPI): Focuses not only on income, but also on education, health, and standard of living.
We define a country as ”underrepresented and economically disadvantaged” if it is classified by the UN as a Least Developed Country (LDC) and appears in the bottom quartile of the Human Development Index (HDI) rankings for its region. To ensure that our research covers a broad range of ’underrepresented’ countries, one from each of the five populated continents (Africa, Europe, North America, South America and Oceania), we select based on the following criteria:
@@ 4.1 Evaluation Methodology
To ensure a controlled evaluation, we keep all hyperparameters constant across models and use an identical prompting strategy. Each model attempts every question three times, generating three independent responses per dataset. The prompt used for the evaluation is provided in Figure A4 (Appendix B 5).
@@ 4.2.1 Model Accuracy Across Cultural Variants
For detailed numerical values corresponding to these accuracy scores, including confidence intervals, refer to Table A2 in (Appendix B 6).
@@ 4.2.2 Performance Gap Analysis
In Haiti for instance, Meta-LLaMA 3.1-8B-Instruct has the highest accuracy drop (4.0%), showing that it struggles with numerical reasoning in Haitian contexts. Most models have a drop of 1-3%, with Claude 3.5-Sonnet dropping only 0.3%. Meta-LLaMA 3.1-8B and 70B have the highest accuracy drop on Moldova( 3.8%), followed by Microsoft Phi-3 Medium. Claude 3.5-Sonnet remains the most stable, with only a 0.4% drop. This shows that models that are heavily optimized for English-centric training data tend to show accuracy drops on Moldova, probably due to insufficient exposure of cultural reasoning patterns in training data. More balanced multilingual training may help address this issue, as shown by models such as Claude 3.5-Sonnet, which demonstrates stronger generalization across various datasets. Similar trends are observed in the other subplots as well. Detailed values can be found in Table A3 and Table A4 in Appendix B 7.
@@
Claude 3.5 consistently outperforms other models across various datasets and is notably robust in mathematical reasoning. It achieved a score of 71.1% on the MATH benchmark and an impressive 91.6% on multilingual math tasks (MGSM) (Anthropic [20]). However, when cultural and linguistic variations are introduced, the performance dynamics changes slightly. Interestingly, Mistral Saba, despite not having a dedicated mathematical reasoning benchmark and not being explicitly tuned for math, handles the Pakistan variant and some other cultural adaptations better. This may be attributed to its training on data from Middle Eastern and South Asian sources. This indicates that even if a model isn’t specifically designed for mathematical reasoning, having a deeper understanding of a region’s cultural context and reasoning patterns can help it perform better on math tasks within that specific cultural framework.
@@
There could be other possible reasons for the accuracy drop of models when the context is shifted. Different languages and cultural terminologies lead to variations in tokenization. If the model’s tokenizer is not well-suited to the culturally adapted language, a single concept might be represented by more tokens, leading to increased complexity and potential for error. The model may also have fewer training examples with the specific terms used in the adapted problems (Petrov et al. [21], Dang et al. [22]). This suggests that models are not entirely language-agnostic for mathematical reasoning performance, even if they are multilingual. Their mathematical reasoning performance could be tied to how efficiently their tokenizers can represent different languages and cultural vocabularies. As shown in Figure A5 (Appendix B 8), adapting names from ‘Amalia, Megan, and Dior’ to ‘Aleskandra, Nicolae, and Albert’ changes the total number of tokens and characters. In the original text, we have 104 tokens and 452 characters, while the Moldovan-adapted version yields 109 tokens and 477 characters. This variation reflects how the tokenizer handles different linguistic structures and could play a role in affecting model performance.
@@
Similarly, mathematical reasoning isn’t always culturally neutral. The way a problem is framed, even with the same numbers, can subtly influence how someone may approach it. Different cultures might have different common-sense assumptions or preferred methods for organizing information (Meng and Liu [23], Tajika [24]). This implies that LLMs aren’t just learning to solve math problems; they’re also learning to solve them in a particular way, based on the dominant problem-solving styles in their training data. Cultural versions might require slightly different reasoning pathways that the model is less familiar with.
Also, LLMs are trained on vast amounts of data. If the pre-training data is heavily biased towards certain cultures or problem-solving styles, the model will naturally perform better in those familiar contexts. The cultural versions of GSM8K introduce scenarios the model hasn’t seen as often (or at all) during pre-training. This underscores the importance of diverse and representative pre-training datasets. A model trained primarily on Western-centric data might struggle with problems framed in the context of other cultures.
Last but not the least, LLMs are prone to hallucination (Banerjee et al. [25]). In culturally adapted versions, models might unintentionally introduce stereotypes or incorrect cultural assumptions into their reasoning process, leading to errors. For example, it might associate a specific region with a certain type of economic activity and introduce that (incorrect) assumption into the calculation.
4.2.3 Statistical Significant Testing
To statistically assess whether LLMs perform differently on culturally adapted math questions compared to the original GSM8K test set, we conduct McNemar tests using one-to-one aligned question pairs. The results reveal a clear pattern: models like LLaMA 3.1-70B, Gemini Flash 2.0, and Mistral Large 2411 consistently exhibit statistically significant performance drops (p $<$ 0.01 across most datasets), with high b-values indicating that these models frequently answered questions correctly on GSM8K but failed on their culturally adapted versions. This supports the rejection of the null hypothesis, suggesting these models struggle more in mathematical reasoning when cultural contexts shift.
In contrast, Claude 3.5 and Mistral Saba generally show no significant performance difference (p $>$ 0.05 in most cases), with balanced b/c values, indicating that their performance is more stable across cultural contexts, and we fail to reject the null hypothesis for these models.
A detailed breakdown of McNemar test statistics, p-values, and b/c counts for each model is provided in Table A5 (Appendix B 9).
@@ 4.2.4 Qualitative Error Analysis
In addition to quantitative evaluations, we conduct a detailed qualitative error analysis to better understand how LLMs handle culturally adapted math word problems. Our analysis reveals three major patterns of reasoning failure across models:
  • Models often struggle with numerical reasoning when using less familiar currency units (e.g., Haitian Gourde, HTG). For instance, some models treat 0.1 HTG as though it were 1 HTG, whereas they correctly handle 0.1 USD—likely because HTG is rarely used in decimals due to inflation and rounding practices, causing the models to miss consistent arithmetic across currency formats.
  • Replacing “wife” with unfamiliar family terms like Jija (Pakistani) or Tambu man (Solomon Islands) often led models to miscalculate. While they correctly process “husband-wife,” they struggle with non-Western family structures—pointing to errors in how they interpret different cultural entities.
@@
These insights highlight how cultural context can introduce variability in reasoning even when the underlying math remains unchanged. Examples and detail of these errors are provided in Figures A6, A7, and A8 (Appendix B 10.1 and B 10.2)
@@ 5 Conclusion
The results show that larger models, such as Anthropic Claude 3.5-Sonnet and GPT-4o, tend to generalize better across these diverse contexts; however, they still experience noticeable performance drops when faced with culturally adapted math problems. In contrast, models like Mistral Saba—despite not being explicitly tuned for math—demonstrate improved performance in regions where they have been exposed to local data. This suggests that cultural familiarity can enhance mathematical problem solving, as models leverage context-specific linguistic and structural cues.
Furthermore, our analysis indicates that changes in cultural context may also pose challenges to the tokenization process. Variations in vocabulary and linguistic structures across different regions can lead to differences in how input text is tokenized, which in turn affects model performance on math problem solving tasks.
In general, our study underscores the importance of considering cultural context when evaluating the mathematical reasoning capabilities of LLMs. It highlights that while universal mathematical principles are at play, factors such as cultural familiarity play significant roles in shaping model performance.
@@ 1 Prompt for Cultural Entities Recognition
2 Evaluation and Manual Correction of Recoganized cultural Entities
To verify that GPT-4o accurately identified only cultural entities, we prepare a 5-shot prompt as shown in the Figure  A3 and use it to verify its own output. The goal is to check whether it correctly identified cultural entities, replaced them with placeholders in the right spots, and ensure that the mathematical logic and structure of the questions remained intact. The prompt can be found in the appendix.
After this step, we conduct human evaluation as well. We manually review symbolic questions and the identify cultural entities to verify that:
  1. 1.
    The mathematical logic of the questions remained unchanged.
  2. 2.
    There were no modifications to numerical values.
@@ 2 Evaluation and Manual Correction of Recoganized cultural Entities
During this process, we encounter a small inconsistency. While GPT-4o correctly identified cultural entities and verified them using our 5-shot prompt, it used inconsistent names for the same entity. For example, it correctly identified a Person name in each question and replaced it with a placeholder. But sometimes, it used name, common name, or Person name in the placeholders. Similarly, there were variations in other entities—like food items, which appeared as food item, common food items, types of food, and cooking items. These technically mean the same thing but have different representative names.
This inconsistency could cause issues in our dictionary creation because we would end up with different keys for the same set of values, unnecessarily inflating the dictionary. To fix this, we modify our 7-shot prompt, run another iteration, and manually review the symbolic version again. Any remaining inconsistent placeholder names are corrected.
@@ 3 Recognized Cultural Entities by GPT-4o
Person name currency
Types of pastries/local deserts currency sign
City name Types of commercial establishments
Types of houses Types of dance
Types of goods merchant purchase Types of common jobs
food items clothing items
Common type of sport Common brand name
cooking item Types of events
Types of beverages Common clothing items
Types of books Types of vehicles
Types of places animal
Recreation activity Types of family events
types of shows Village names
School subject cultural event
Types of games Types of flowers
family member recreation places
Types of musical compositions profession
Types of classes holiday
company names Types of teacher
restaurant name cultural landmark
Mythical character online shopping platforms
Types of entertainment places cultural dance style
Government body Types of scents
Cultural songs school name
common places Types of tea
appliances newspaper names
religious place Language
school subject
Table A1: Cultural Entities

### v2

@@ 4 Country Selection
We define a country as “underrepresented and economically disadvantaged” if it is classified by the UN as a Least Developed Country (LDC) and appears in the bottom quartile of the Human Development Index (HDI) rankings for its region. To ensure that our research covers a broad range of ‘underrepresented’ countries, one from each of the five populated continents (Africa, Europe, North America, South America and Oceania), we select based on the following criteria:
@@ 5.1 Evaluation Methodology
To ensure a controlled evaluation, we keep all hyperparameters at their default values constant across models and use an identical prompting strategy. Specifically, no temperature parameter was explicitly set, ensuring that all models were evaluated under identical API default configurations rather than artificially tuned settings. Each model attempts every question three times, generating three independent responses per dataset. The prompt used for the evaluation is provided in Figure A4 (Appendix B).
@@ 5.2.1 Model Accuracy Across Cultural Variants
For detailed numerical values corresponding to these accuracy scores, including confidence intervals, refer to Table A3 in (Appendix B).
@@ 5.2.2 Performance Gap Analysis
In Haiti for instance, Meta-LLaMA 3.1-8B-Instruct has the highest accuracy drop (4.0%), showing that it struggles with numerical reasoning in Haitian contexts. Most models have a drop of 1-3%, with Claude 3.5-Sonnet dropping only 0.3%. Meta-LLaMA 3.1-8B and 70B have the highest accuracy drop on Moldova( 3.8%), followed by Microsoft Phi-3 Medium. Claude 3.5-Sonnet remains the most stable, with only a 0.4% drop. This shows that models that are heavily optimized for English-centric training data tend to show accuracy drops on Moldova, probably due to insufficient exposure of cultural reasoning patterns in training data. While models such as Claude 3.5-Sonnet demonstrate stronger generalization across cultural variants, the specific factors driving this robustness, whether model scale, training data diversity, or stronger mathematical reasoning capabilities, cannot be extracted from our experiments. Detailed values can be found in Table A4 and Table A5 (Appendix B).
@@
Claude 3.5 consistently outperforms other models across various datasets and is notably robust in mathematical reasoning. It achieved a score of 71.1% on the MATH benchmark and an impressive 91.6% on multilingual math tasks (MGSM) (Anthropic,, 2024). However, when cultural and linguistic variations are introduced, the performance dynamics changes slightly. Interestingly, Mistral Saba, despite not having a dedicated mathematical reasoning benchmark and not being explicitly tuned for math, handles the Pakistan variant and some other cultural adaptations better. This may be attributed to its training on data from Middle Eastern and South Asian sources. This indicates that even if a model isn’t specifically designed for mathematical reasoning, having a deeper understanding of a region’s cultural context and reasoning patterns can help it perform better on math tasks within that specific cultural framework.
@@
There could be other possible reasons for the accuracy drop of models when the context is shifted. Different languages and cultural terminologies may lead to variations in tokenization. If the model’s tokenizer is not well-suited to the culturally adapted language, a single concept might be represented by more tokens, potentially increasing input complexity. The model may also have fewer training examples with the specific terms used in the adapted problems (Petrov et al.,, 2023), (Dang et al.,, 2024). This suggests that models are not entirely language-agnostic for mathematical reasoning performance, even if they are multilingual. Their mathematical reasoning performance may correlate with how efficiently their tokenizers can represent different languages and cultural vocabularies, though establishing a direct causal link falls outside the scope of this black-box evaluation. As shown in Figure A5 (Appendix B), adapting names from ‘Amalia, Megan, and Dior’ to ‘Aleskandra, Nicolae, and Albert’ changes the total number of tokens and characters. In the original text, we have 104 tokens and 452 characters, while the Moldovan-adapted version yields 109 tokens and 477 characters. This variation reflects how the tokenizer handles different linguistic structures and represents one potential contributing factor among several possible explanations for the observed performance differences.
@@
Similarly, mathematical reasoning isn’t always culturally neutral. The way a problem is framed, even with the same numbers, can subtly influence how someone may approach it. Different cultures might have different common-sense assumptions or preferred methods for organizing information (Meng and Liu,, 2022), (Tajika,, 2004). This implies that LLMs aren’t just learning to solve math problems; they’re also learning to solve them in a particular way, based on the dominant problem-solving styles in their training data. Cultural versions might require slightly different reasoning pathways that the model is less familiar with. If the pre-training data is heavily biased towards certain cultures or problem-solving styles, the model will naturally perform better in those familiar contexts. This underscores the importance of diverse and representative pre-training datasets.
5.2.3 Statistical Significance Testing
To statistically assess whether LLMs perform differently on culturally adapted math questions compared to the original GSM8K test set, we conduct McNemar tests using one-to-one aligned question pairs. The results reveal a clear pattern: models like LLaMA 3.1-70B, Gemini Flash 2.0, and Mistral Large 2411 consistently exhibit statistically significant performance drops (p << 0.01 across most datasets), with high b-values indicating that these models frequently answered questions correctly on GSM8K but failed on their culturally adapted versions. This supports the rejection of the null hypothesis, suggesting these models struggle more in mathematical reasoning when cultural contexts shift.
In contrast, Claude 3.5 and Mistral Saba generally show no significant performance difference (p >> 0.05 in most cases), with balanced b/c values, indicating that their performance is more stable across cultural contexts, and we fail to reject the null hypothesis for these models.
A detailed breakdown of McNemar test statistics, p-values, and b/c counts for each model is provided in Table A6 (Appendix B).
@@ 5.2.4 Qualitative Error Analysis
We conduct a detailed qualitative error analysis to better understand how LLMs handle culturally adapted math word problems. The qualitative error analysis was conducted through systematic manual inspection of model outputs across all 14 models and all six cultural variants. Full methodology and examples are provided in Appendix B Section 11. Our analysis reveals three major patterns of reasoning failure across models:
  • Models often struggle with numerical reasoning when using less familiar currency units (e.g., Haitian Gourde, HTG). For instance, some models treat 0.1 HTG as though it were 1 HTG, whereas they correctly handle 0.1 USD, likely because HTG is rarely used in decimals due to inflation and rounding practices, causing the models to miss consistent arithmetic across currency formats.
  • Replacing “wife” with unfamiliar family terms like Jija (Pakistani) or Tambu man (Solomon Islands) often led models to miscalculate. While they correctly process “husband-wife,” they struggle with non-Western family structures, pointing to errors in how they interpret different cultural entities.
@@
These insights highlight how cultural context can introduce variability in reasoning even when the underlying math remains unchanged. Examples and detail of these errors are provided in Figures A6, A7, and A11 (Appendix B)
5.2.5 Quantitative Error Analysis
While Section  5.2.4 illustrates specific failure patterns through representative examples, we systematically quantify the prevalence of different error types across all models and cultural variants. We analyze 18,887 error instances identified using the strict consistency criterion described in Section  5.1, where at least one of the three model runs produced an incorrect answer. These errors are aggregated across all 14 evaluated models and all seven datasets, the original GSM8K and its six cultural variants, and include all incorrect responses regardless of whether the model answered the original GSM8K question correctly or not. Each error is categorized into six types, as shown in the Table  2, using an LLM-assisted annotation process, which is validated through manual review of 50 randomly sampled cases (detailed methodology in Appendix B Section  12).
Error Category Count Percentage (%)
Mathematical Reasoning error 10,331 54.70
Calculation error 6,523 34.54
Relationship misunderstanding 851 4.51
Unit / currency error 746 3.95
Problem misinterpretation 292 1.55
Other / unclear 144 0.76
Total 18,887 100.00
Table 2: Overall distribution of error types across all models and cultural variants.
Table  2 presents the overall distribution of error types. The analysis reveals that mathematical reasoning errors constitute the majority of failures (54.7%), followed by calculation errors (34.5%). Notably, culturally-specific errors, relationship misunderstandings (4.5%) and unit/currency errors (4.0%), comprise a relatively small proportion of total errors. This distribution quantifies the error patterns identified in Section  5.2.4, revealing that while explicitly culturally-specific errors (relationship and currency) represent 8.5% of failures, the broader impact of cultural adaptation on model reasoning is reflected across all error categories. Representative examples of each error type are provided in Appendix B Section 11. A complementary analysis examining which cultural entity categories are disproportionately associated with incorrect responses is provided in Appendix B Section 12.3.
@@ 6 Conclusion
The results show that larger models, such as Anthropic Claude 3.5-Sonnet and GPT-4o, tend to generalize better across these diverse contexts; however, they still experience noticeable performance drops when faced with culturally adapted math problems. In contrast, models like Mistral Saba, despite not being explicitly tuned for math, demonstrate improved performance in regions where they have been exposed to local data, suggesting that cultural familiarity can enhance mathematical problem solving. Furthermore, our analysis indicates that changes in cultural context may also pose challenges to the tokenisation process, as variations in vocabulary and linguistic structures across different regions can lead to differences in how input text is tokenised.
In general, our study underscores the importance of considering cultural context when evaluating the mathematical reasoning capabilities of LLMs. Beyond evaluation, the culturally adapted benchmarks introduced in this study could serve as a foundation for future work in training more culturally robust LLMs, for instance, through targeted data augmentation strategies that expose models to underrepresented cultural contexts during fine-tuning. The dataset creation pipeline presented here is also replicable and extensible, offering a template for developing similar benchmarks across additional cultures and languages.
@@ 1 Prompt for Cultural Entities Recognition
2 Validation of Cultural Entity Recognition Quality
2.1 Sampling Methodology
We randomly sampled 130 questions from the 1,198 culturally adaptable questions in our dataset, representing approximately 11% of the total. This sample size provides sufficient statistical power for quality assessment while remaining feasible for detailed manual review. The random sampling was performed without stratification to ensure an unbiased representation of the dataset’s diversity in terms of question complexity, number of entities per question, and entity types.
2.2 Annotation Guidelines and Evaluation Criteria
Two independent human annotators (authors of this paper with expertise in the dataset creation process) were tasked with evaluating GPT-4o’s entity recognition output. For each of the 130 sampled questions, annotators were instructed to answer a binary question: “Did GPT-4o correctly identify all cultural entities in this question without omissions or errors?” The evaluation criteria were defined as follows:
  • Yes (Correct): GPT-4o identified all cultural entities present in the question with no omissions and no incorrect identifications.
  • No (Incorrect): GPT-4o either (1) missed at least one cultural entity, (2) incorrectly identified a non-entity as a cultural entity, or (3) misclassified the entity type.
This strict, question-level evaluation approach ensures that only questions with perfect entity recognition are marked as correct, providing a conservative estimate of GPT-4o’s performance.
2.3 Validation Results
The two human annotators as well as GPT-4o, evaluated all 130 sampled questions independently. The results are summarized in Table  A1:
Evaluator Correct (Yes) Incorrect (No) Accuracy (%)
Human Annotator 1 96 34 73.8
Human Annotator 2 103 27 79.2
GPT-4o (Self-evaluation) 86 44 66.2
Table A1: Validation results for cultural entity recognition on 130 sampled questions.
The validation results show that the two human annotators achieved question-level accuracy of 73.8% and 79.2%, respectively, indicating that GPT-4o successfully identified all cultural entities without errors in approximately three out of four questions. Additionally, we employed GPT-4o itself as a third evaluator to assess its own output, which yielded 86 correct identifications (66.2% accuracy), suggesting it applied stricter criteria when assessing its own output.
2.4 Correction and Iterative Refinement
All questions identified as incorrect by any of the evaluators were manually reviewed to identify the specific errors in GPT-4o’s entity recognition. Errors fell into two categories: placeholder naming inconsistencies, addressed through the prompt refinement process described in Appendix A Section  3, and broader entity recognition errors including missed entities and incorrect entity type classifications, addressed through the multi-layer correction process described below.
To address broader entity recognition errors across the full dataset, we employed GPT-4o as an automated evaluator on all 1,198 culturally adaptable questions. Unlike its role as a third evaluator in the 130-question validation study, here GPT-4o was explicitly instructed with the specific error patterns identified during human validation, including missed entities and incorrect entity type classifications, and applied systematically across every question in the dataset.
In parallel, we refined the entity recognition 7-shot prompt iteratively based on the error patterns discovered during both the human validation and GPT-4o automated evaluation. Each refined version of the prompt was reapplied to the complete dataset, with subsequent evaluation cycles used to verify improvement. This iterative prompt refinement addressed both placeholder naming inconsistencies, detailed in Appendix A Section  3, and broader entity recognition errors across the full dataset.
Finally, manual human inspection was conducted across the full dataset by the authors, drawing on our cultural familiarity with the target regions, particularly Pakistan, to identify and correct culturally sensitive errors that automated evaluation could not reliably assess. This was especially important for entity types requiring cultural understanding, such as regional family terms, local food items, and culturally specific place names. The base evaluation prompt structure used to guide GPT-4o as a judge throughout this process is provided in Appendix A Figure  A2, with targeted refinements applied iteratively based on errors discovered in each correction cycle.
The validation study confirms that while GPT-4o’s initial entity recognition achieved reasonable accuracy (73-79% as judged by human annotators), manual oversight and iterative correction, the multi-layer correction process, combining automated evaluation, iterative prompt refinement, and manual human inspection, was essential to ensure the quality and reliability of our culturally adapted datasets.
3 Iterative Correction of Placeholder Naming Inconsistencies
@@
During the validation study described in Appendix A Section 2, we employed a 5-shot evaluation prompt (Figure  A2) to evaluate GPT-4o’s entity recognition output. This evaluation process revealed a naming inconsistency issue in GPT-4o’s placeholder assignments. While GPT-4o correctly identified cultural entities, it used inconsistent placeholder names for the same entity type across different questions. For example, person names were variously labeled as name, common name, or Person name. Similarly, food-related entities appeared as food item, common food items, types of food, and cooking items.
This inconsistency posed a problem for dictionary creation, as different placeholder names for the same entity type would create redundant keys with identical value sets. To address this issue, we modified the 7-shot prompt (Figure A1) to enforce consistent placeholder naming conventions and conducted additional iterations of entity recognition.
@@ 4 Recognized Cultural Entities by GPT-4o
Category Entity Type %
People and Roles Person name 87.1%
Family member 4.9%
Types of common jobs 2.0%
Profession 1.0%
Places and Locations Types of commercial establishments 5.7%
City name 1.8%
Types of places 1.0%
Restaurant name 0.8%
School name 0.7%
Types of entertainment places 0.5%
Types of houses 0.4%
Common places 0.2%
Religious place 0.2%
Village names 0.1%
Recreation places 0.1%
Cultural landmark 0.1%
Government body 0.1%
Food and Drink Food items 11.2%
Cooking item 2.1%
Types of beverages 1.2%
Types of pastries/local desserts 0.1%
Types of tea 0.1%
Education School subject 0.4%
Types of classes 0.3%
Types of teacher 0.3%
Clothing and Appearance Clothing items 1.5%
Common clothing items 1.5%
Types of flowers 0.4%
Types of scents 0.1%
Category Entity Type %
Culture and Arts Types of events 2.2%
Cultural event 1.2%
Types of family events 1.0%
Holiday 0.5%
Types of books 0.3%
Mythical character 0.3%
Types of shows 0.2%
Types of musical compositions 0.2%
Recreation activity 0.2%
Types of dance 0.1%
Cultural dance style 0.1%
Cultural songs 0.1%
Types of games 0.1%
Commerce and Economy Currency sign 31.8%
Currency 20.7%
Company names 0.5%
Common brand name 0.4%
Types of goods merchant purchase 0.1%
Online shopping platforms 0.1%
Appliances 0.1%
Other Animal 2.3%
Common type of sport 1.3%
Types of vehicles 0.2%
Language 0.1%
Newspaper names 0.1%
Table A2: Cultural Entity Types by Category with Occurrence Percentages
Table  A2 presents the 54 cultural entity types identified across the 1,198 culturally adaptable questions, organised into eight broad categories. The percentage indicates the proportion of questions in which each entity type appears at least once. Person name is the most frequently occurring entity type, present in 87.1% of questions, followed by currency sign (31.8%) and currency (20.7%), reflecting the nature of GSM8K as a math word problem dataset involving people and monetary values. Most other entity types appear in fewer than 5% of questions.